OpenStudy (hdrager):

solve ln(y-1)-ln 2 = x + ln x

OpenStudy (hdrager):

solve for y

OpenStudy (irishboy123):

\( \ln(y-1)- \ln 2 = x + \ln x\) some starting hints \( \ln \dfrac{y-1}{ 2} = x + \ln x\) \(\ln \dfrac{y-1}{ 2} = \ln e^x + \ln x\) 2 more steps!

OpenStudy (hdrager):

wait why was x on the right side of the equation put int exponential form? @IrishBoy123

OpenStudy (irishboy123):

because you can then say that \( \ln e^x + \ln x\) \(= \ln x e^x \)

OpenStudy (irishboy123):

make sense?

OpenStudy (hdrager):

so do you then multiply by 2 and then add 1 on each side?

OpenStudy (irishboy123):

well, you've now got this \(\ln \dfrac{y-1}{ 2} = \ln x e^x\) so you can "un-log" it, if that makes sense

OpenStudy (hdrager):

how do you do that? do you divide both sides by ln?

OpenStudy (irishboy123):

nah! just "un-function" them if \( \ln P = \ln Q\) then \(P = Q\)

OpenStudy (hdrager):

ohhhhhh okay so they just automatically cancel out one another

OpenStudy (irishboy123):

yes!

OpenStudy (irishboy123):

So \( \dfrac{y-1}{ 2} = x e^x\)

OpenStudy (irishboy123):

still needs to be finished off

OpenStudy (hdrager):

y = 2xe^x+1

OpenStudy (irishboy123):

\(\Huge \checkmark\)

OpenStudy (hdrager):

YES thanks so much!!

OpenStudy (irishboy123):

mp!